# Random Utility with Aggregation

### v2

@@ Appendix A Proofs
[…]
$=\rho(D\cup E,z),$
A.4 Proof of Theorem 3.6 and Remark 3.10
A.4.1 Proof of Theorem 3.6
First we prove that the set of RU rational choice functions coincides with $\text{co.}\{\rho^{\succ}_{\mathscr{E}}\mid\succ\in\mathscr{L}(\mathscr{A})\text{ and }\mathscr{E}\subseteq\mathscr{D}\}$.
Fix $\succ\in\mathscr{L}(\mathscr{A}_{N}\cup X(a_{0}))$. For each $A\in\mathscr{D}$, fix $(S_{a})_{a\in A}\in\prod_{a\in A}(2^{X(a)}\setminus\{\emptyset\})$. For each collection $((S_{a})_{a\in A})_{A\in\mathscr{D}}$, we define $\rho^{\succ}_{((S_{a})_{a\in A})_{A\in\mathscr{D}}}$ as follows. To make the notation simple, we write $((S_{a})_{a\in A})_{A\in\mathscr{D}}$ as $(S_{A})_{A\in\mathscr{D}}$. For each $b\in B\in\mathscr{D}$,
$\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(B,b)=1\Big(\exists x\in S_{b}\forall y\in\cup_{a\in B\setminus b}S_{a}[x\succ y]\Big).$
Step 1: For each $\succ\in\mathscr{L}(\mathscr{A}_{N}\cup X(a_{0}))$ and each collection $(S_{A})_{A\in\mathscr{D}}$, there exists $\succ^{\prime}\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\subseteq\mathscr{D}$ such that $\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}=\rho^{\succ^{\prime}}_{\mathscr{E}}$.
Proof.
Fix each $\succ\in\mathscr{L}(\mathscr{A}_{N}\cup X(a_{0}))$ and each collection $(S_{A})_{A\in\mathscr{D}}$, define
${\cal F}=\Bigg\{D\in\mathscr{D}\ \Bigg|\ \forall y\in\bigcup_{b\in D\setminus a_{0}}X(b)\ [x^{*}\succ y]\Bigg\},$
where $x^{*}=\max_{S_{a_{0}}}\succ$. That is, $\mathcal{F}$ consists of all choice sets $D$ such that the $\succ$-best element lies in $S_{a_{0}}$ and is preferred to every element contained in $\bigcup_{b\in D\setminus a_{0}}X(b)$. Define $\mathscr{E}=\mathscr{D}\setminus{\cal F}$.
Define $\succ^{\prime}\in\mathscr{L}(\mathscr{A}_{N})$ such that $a\succ^{\prime}b$ if and only if $x\succ y$ for any $a,b\in\mathscr{A}_{N}$, where $\{x\}=X(a)$ and $\{y\}=X(b)$. For any $D\in\mathscr{D}$, if $D\in\mathscr{E}^{c}\equiv{\cal F}$, the maximum elements of $\succ$ over $D$ belongs to $X(a_{0})$, thus $\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(D,\cdot)=\rho^{\succ^{\prime}}_{\mathscr{E}}(D,\cdot)$. Moreover, if $D\in\mathscr{E}$, the maximum elements of $\succ$ and $\succ^{\prime}$ over $D$ coincide; thus $\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(D,\cdot)=\rho^{\succ^{\prime}}_{\mathscr{E}}(D,\cdot)$. ∎
Given Step 1, it suffices to show that any RU-rational $\rho$ can be expressed as a convex combination of terms of the form $\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}$. Assume that $\rho$ is rationalized by $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$. We first establish this in the degenerate case where the rationalizing measure $\mu_{{\scriptscriptstyle\mathscr{X}}}$ is a point mass in Step 2.191919The general case is proved in Step 3 In particular, we show that $\rho$ can be represented as a convex combination of $\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}$, with weights $\alpha((S_{A})_{A\in\mathscr{D}})$ defined by
$\alpha\big((S_{A})_{A\in\mathscr{D}}\big)\;\equiv\;\prod_{A\in\mathscr{D}}\lambda_{A}(S_{A}).$
Since each $\lambda_{A}$ is a probability distribution over $S_{A}$, the product measure $\alpha$ is also a probability distribution over the entire collection $(S_{A})_{A\in\mathscr{D}}$.
Step 2: Suppose that $\rho$ is RU-rationalized by $(\delta_{\succ},\lambda)$. For each $(B,b)$ such that $b\in B\in\mathscr{D}$,
[…]
Proof.
For each $(B,b)$ such that $b\in B\in\mathscr{D}$, the left hand side equals
$\rho(B,b)=\sum_{(S_{b})_{b\in B}\in\prod_{b\in B}(2^{X(b)}\setminus\{\emptyset\})}\lambda_{B}((S_{b})_{b\in B})1(\exists x\in S_{b}\forall y\in\cup_{a\in B\setminus b}S_{a}[x\succ y]).$
As for the right hand side
$\sum_{(S_{A})_{A\in\mathscr{D}}\in\prod_{A\in\mathscr{D}}\text{ supp }\lambda_{A}}\alpha((S_{A})_{A\in\mathscr{D}})\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(B,b)$
$=\sum_{(S_{A})_{A\in\mathscr{D}}\in\prod_{A\in\mathscr{D}}\text{ supp }\lambda_{A}}\Big(\prod_{A\in\mathscr{D}}\lambda_{A}(S_{A})\Big)\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(B,b)$
$=\sum_{S_{B}\in\text{ supp }\lambda_{B}}\lambda_{B}(S_{B})\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(B,b)\times\sum_{(S_{A})_{A\in\mathscr{D}\setminus B}\in\prod_{A\in\mathscr{D}\setminus B}\text{ supp }\lambda_{A}}\Big(\prod_{A\in\mathscr{D}\setminus B}\lambda_{A}(S_{A})\Big)$
$\quad(\because\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(B,b)\text{ depends only on }S_{B})$
$=\sum_{S_{B}\in\text{ supp }\lambda_{B}}\lambda_{B}(S_{B})\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}(B,b)\quad(\because\text{ the second term is $1$})$
$=\sum_{(S_{b})_{b\in B}\in\prod_{b\in B}(2^{X(b)}\setminus\{\emptyset\})}\lambda_{B}(S_{B})1(\exists x\in S_{b}\forall y\in\cup_{a\in B\setminus b}S_{a}[x\succ y]).$
Given Step 2, the general case follows. To see this suppose that $\rho$ is RU-rationalized by $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$. Observe that for any $(A,a)$ such that $a\in A\in\mathscr{D}$,
$\rho(A,a)=\sum_{\succ\in\mathscr{L}(\mathscr{A}_{N}\cup X(a_{0}))}\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ)\underbrace{\sum_{(S_{a})_{a\in A}\in\prod_{a\in A}(2^{X(a)}\setminus\{\emptyset\})}\lambda_{A}((S_{a})_{a\in A})1(\exists x\in S_{a}\forall y\in\cup_{b\in A\setminus a}S_{b}[x\succ y])}_{(*)}.$ (23)
Step 2 shows that each term (*) in equation (23) can be written as a convex combination of elements of the form $\rho^{\succ}_{(S_{A})_{A\in\mathscr{D}}}$. Hence, the entire right-hand side of (23) is itself a convex combination of such terms with $\mu_{{\scriptscriptstyle\mathscr{X}}}$.
We have just shown that the set of RU-rational choice functions is a subset of $\text{co.}\{\rho^{\succ}_{\mathscr{E}}|\succ\in\mathscr{L}(\mathscr{A})\text{ and }\mathscr{E}\subseteq 2^{\mathscr{A}}\}$. [^20] This holds regardless of the size of $X(a_{0})$.
Now we show the other inclusion. By Theorem 3.1, it can be shown that the set of RU-rational choice functions is convex since . Since $\rho^{\succ}_{\mathscr{E}}$ is RU-rational for all $\succ\in\mathscr{L}(\mathscr{A})$ and $\mathscr{E}\in\mathscr{D}$ (see Remark 3.5), thus its convex combination is also RU rational; thus the set $\text{co.}\{\rho^{\succ}_{\mathscr{E}}|\succ\in\mathscr{L}(\mathscr{A})\text{ and }\mathscr{E}\subseteq 2^{\mathscr{A}}\}$ is contained by the set of RU-rational choice functions. Therefore, the RU rational set coincides with $\text{co.}\{\rho^{\succ}_{\mathscr{E}}|\succ\in\mathscr{L}(\mathscr{A})\text{ and }\mathscr{E}\subseteq 2^{\mathscr{A}}\}$.
[…]
A.4.2 Proof of Remark 3.10
We obtain the lower bound of the number of vertices. Fix $\mathscr{E}$ that contains all subsets of $\mathscr{A}$ that contains $a_{0}$. Note that there are $2^{(2^{|\mathscr{A}_{N}|}-{|\mathscr{A}_{N}|\choose 2}-1)}$ distinct such collection $\mathscr{E}$. Let $\cal C$} be the collection of such $\mathscr{E}$s. Let $\mathscr{L}^{\prime}$ be the set of linear orders on $\mathscr{A}$ such that $a_{0}$ is the worst alternative. There are $|\mathscr{A}_{N}|!$ distinct such linear orders.
$|\mathscr{A}_{N}|!\times 2^{(2^{|\mathscr{A}_{N}|}-{|\mathscr{A}_{N}|\choose 2}-1)}.$
Finally we show $\rho^{\succ}_{\mathscr{E}}\neq\rho^{\succ^{\prime}}_{\mathscr{E}^{\prime}}$ if $(\succ,\mathscr{E})\neq(\succ^{\prime},\mathscr{E}^{\prime})$. To see this suppose by way of contradiction that $\rho^{\succ}_{\mathscr{E}}=\rho^{\succ^{\prime}}_{\mathscr{E}^{\prime}}$. Since $\succ,\succ^{\prime}\in\mathscr{L}^{\prime}$, the worst elements of $\succ$ and $\succ^{\prime}$ are $a_{0}$. Moreover, both $\mathscr{E}$ and $\mathscr{E}^{\prime}$ contain the sets of the form $\{a,b,a_{0}\}$ for all $\{a,b\}\subseteq\mathscr{A}_{N}$. Thus, $\rho^{\succ}_{\mathscr{E}}=\rho^{\succ^{\prime}}_{\mathscr{E}^{\prime}}$ implies that $\succ$ and $\succ^{\prime}$ coincide binary comparison of any $a,b\in\mathscr{A}_{N}$. It follows that $\succ=\succ^{\prime}$, which in turn implies $\mathscr{E}=\mathscr{E}^{\prime}$ given $\rho^{\succ}_{\mathscr{E}}=\rho^{\succ^{\prime}}_{\mathscr{E}^{\prime}}$ and the fact that $\mathscr{E}$ contains all sets containing $a_{0}$. [^21] There are many redundant vertices. For example, for any $\mathscr{E}$, fix two linear orders $\succ$ and $\succ^{\prime}$ such that $a_{1}\succ\cdots\succ a_{|\mathscr{A}_{N}|}\succ a_{0}$ and $a_{1}\succ^{\prime}\cdots\succ^{\prime}a_{|\mathscr{A}_{N}|-1}\succ^{\prime}a_{0}\succ^{\prime}a_{|\mathscr{A}_{N}|}$. Then we have $\rho^{\succ}_{\mathscr{E}}=\rho^{\succ^{\prime}}_{\mathscr{E}}$.
A.5 Proofs for Section 4
To prove Proposition 4.2, we first provide a lemma:
Lemma A.1.
Let $\rho$ be a stochastic choice function.
  • (a)
    Suppose that $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ rationalizes $\rho$ and $\mu_{{\scriptscriptstyle\mathscr{X}}}$ is non-overlapping. Then for any composition distribution $\lambda^{\prime}$, $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda^{\prime})$ rationalizes $\rho$.
  • (b)
    Suppose that there is no non-overlapping $\mu_{{\scriptscriptstyle\mathscr{X}}}$ such that $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ rationalizes $\rho$ for some $\lambda$. Then, for any composition distribution $\lambda^{\prime}$, there does not exist non-overlapping $\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime}$ such that $(\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime},\lambda^{\prime})$ rationalizes $\rho$.
Proof.
Statement (a) follows from the fact that if $\mu_{{\scriptscriptstyle\mathscr{X}}}$ is non-overlapping then for all $(S_{a})_{a\in A}\in\prod_{a\in A}(2^{X(a)}\setminus\{\emptyset\})$ and $(S^{\prime}_{a})_{a\in A}\in\prod_{a\in A}(2^{X(a)}\setminus\{\emptyset\})$ we have that $\rho_{\scriptscriptstyle\mathscr{X}}(\cup_{a\in A}S_{a},S_{b})=\rho_{\scriptscriptstyle\mathscr{X}}(\cup_{a\in A}S^{\prime}_{a},S^{\prime}_{b})$ for all for all $b\in A$ where $\rho_{\scriptscriptstyle\mathscr{X}}$ is the underlying stochastic choice function on $\mathscr{X}$ rationalized by $\mu\in\Delta(\mathscr{L}(\mathscr{A}_{N}\cup X(a_{0})))$. Fix $(S^{*}_{a})_{a\in A}$. Then for all $b\in A$, we have
$=\rho_{\scriptscriptstyle\mathscr{X}}(\cup_{a\in A}S^{*}_{a},S^{*}_{b})$
To prove statement (b), assume there is no non-overlapping $\mu_{{\scriptscriptstyle\mathscr{X}}}$ that RU rationalizes $\rho$. By way of contradiction assume that there exist $\lambda^{\prime}$ and non-overlapping $\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime}$ such that $(\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime},\lambda^{\prime})$ that rationalizes $\rho$. Given this, statement (a) implies a contradiction. ∎
A.5.1 Proof of Proposition 4.2
Only if direction: It suffices to show this for the case where $\mu_{{\scriptscriptstyle\mathscr{X}}}$ is deterministic. Suppose $\mu_{{\scriptscriptstyle\mathscr{X}}}$ is non-overlapping and $\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ)=1$ for some $\succ\in\mathscr{L}(\mathscr{X})$ and fix $\lambda$. Label the elements of $\mathscr{A}$ as $a_{1},a_{2},...a_{|\mathscr{A}|}$ in such a way that if $i<j$ then $x\succ y$ for all $x\in X(a_{i})$ and $y\in X(a_{j})$. This labeling is possible by the almost non-overlapping condition.
Let $\succ_{\mathscr{A}}$ be defined such that
$a_{1}\ \succ_{\mathscr{A}}\ a_{2}\ \succ_{\mathscr{A}}\ \cdots\succ_{\mathscr{A}}a_{|\mathscr{A}|}.$
It is easy to see that if $D\subseteq\mathscr{A}$ contains any element of $\mathscr{A}$ then $\rho$ puts probability $1$ on the $a_{i}$ with the largest index. Thus, $\mu_{{\scriptscriptstyle\mathscr{A}}}$ rationalizes $\rho$.
If direction: There exists $\mu_{{\scriptscriptstyle\mathscr{A}}}\in\Delta({\mathscr{L}}({\mathscr{A}}))$ that rationalizes $\rho$. For each $\succ\in\text{supp }\mu_{{\scriptscriptstyle\mathscr{A}}}$ consider a ranking $\succ^{\prime}\in{\mathcal{L}}(\mathscr{X})$ that extends $\succ$ by
$x\succ^{\prime}y\text{ if }x\in X(a)\text{ and }y\in X(b)\text{ for some }a\ \succ\ b$
and fill the order among $X(a)$ arbitrarily for all $a$; then define $\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ^{\prime})=\mu_{{\scriptscriptstyle\mathscr{A}}}(\succ)$. Also define $\lambda$ arbitrarily. Since $\mu_{{\scriptscriptstyle\mathscr{X}}}$ is non-overlapping, it follows from statement (b) of Lemma A.1 that $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ rationalizes $\rho$.
A.5.2 Proof of Proposition 4.5
If direction: If $\rho$ is rationalizable by a random utility model $\mu_{{\scriptscriptstyle\mathscr{A}}}$ over $\mathscr{A}$, then as we showed in Proposition 4.2 that there exists $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ that rationalizes $\rho$ and $\mu_{{\scriptscriptstyle\mathscr{X}}}$ is non-overlapping. Moreover, as proved in Lemma A.1 (a), for non-overlapping representation any $\lambda$ works; so we can choose an independent $\lambda$ as desired.
Only if direction: Suppose that $\rho$ is RU rationalizable with $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$, where $\lambda$ is independent. Let $\rho_{\scriptscriptstyle\mathscr{X}}$ be the random utility function associated with $\mu_{{\scriptscriptstyle\mathscr{X}}}$.
When $\lambda$ is independent, representation (3) in Definition 2.5 can be simplified as in the following lemma.
Lemma A.2.
Let $\rho$ be a stochastic choice function. If $\rho$ is rationalized by $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ where $\lambda$ is menu-independent, we have for all $A\subseteq{\mathscr{A}}$ and $a\in A$
$\rho(A,a)=\sum_{(S_{b})_{b\in{\mathscr{A}}}}\lambda((S_{b})_{b\in\mathscr{A}})\rho_{\mathscr{X}}(\cup_{b\in A}S_{b},S_{a}),$
where $\rho_{\mathscr{X}}(\cup_{b\in A}S_{b},S_{a})=\mu_{{\scriptscriptstyle\mathscr{X}}}(\succ\in\mathscr{L}(\mathscr{X})|\exists x\in S_{a},\ \forall y\in\cup_{b\in A\setminus a}S_{b},\ x\succ y)$
Lemma A.2 means that the aggregated stochastic choice function $\rho$ is a convex combination of $\rho_{\mathscr{X}}$ for each composition of the aggregated categories.
Note that, by definition, $\mu_{(S_{a})_{a\in\mathscr{A}}}$ RU rationalizes $\rho_{(S_{a})_{a\in\mathscr{A}}}$, that is for any $(A,a)$ such that $a\in A\subseteq\mathscr{A}$, $\rho_{(S_{a})_{a\in\mathscr{A}}}(A,a)=\mu_{(S_{a})_{a\in\mathscr{A}}}(\succ^{\prime}\in{\mathscr{L}}({\mathscr{A}})|a\succ^{\prime}b,\text{ for all }b\in A\setminus a)$. Now let $\mu_{{\scriptscriptstyle\mathscr{A}}}=\sum_{(S_{a})_{a\in{\mathscr{A}}}}\lambda((S_{a})_{a\in\mathscr{A}})\mu_{(S_{a})_{a\in\mathscr{A}}}$. Then
[…]
Thus $\mu_{{\scriptscriptstyle\mathscr{A}}}$ rationalizes $\rho$.
A.5.3 Proof of Lemma A.2
Fix $A\subseteq\mathscr{A}$ and $a\in A$. Then,
$=\sum_{\left(S_{b}\right)_{b\in A}\in\prod_{b\in A}\left(2^{X(b)}\setminus\{\emptyset\}\right)}\lambda\Big(\prod_{b\in A}(S_{b})\times\prod_{b\in\mathscr{A}\setminus A}(2^{X(b)}\setminus\{\emptyset\})\Big)\rho_{\scriptscriptstyle\mathscr{X}}(\cup_{b\in A}S_{b},S_{a})$
$=\sum_{\left(S_{b}\right)_{b\in A}\in\prod_{b\in A}\left(2^{X(b)}\setminus\{\emptyset\}\right)}\sum_{\left(S_{c}\right)_{a\in B^{c}}\in\prod_{c\in B^{c}}\left(2^{X(c)}\setminus\{\emptyset\}\right)}\lambda((S_{a})_{a\in B},(S_{c})_{c\in B^{c}})\rho_{\scriptscriptstyle\mathscr{X}}(\cup_{b\in A}S_{b},S_{a})$
$=\sum_{\left(S_{b}\right)_{b\in\mathscr{A}}\in\prod_{b\in\mathscr{A}}(2^{X(b)}\setminus\{\emptyset\})}\lambda\left((S_{b})_{b\in\mathscr{A}}\right)\rho_{\scriptscriptstyle\mathscr{X}}(\cup_{b\in B}S_{b},S_{a}).$
Appendix B Online Appendix: Supplementary Simulation Results–Bias and Distance
B.1 Estimation Bias
B.1.1 Effect of composition distributions $\lambda$

### v3

@@ D.2 Technical remarks
$\{x,a_{0}\}\notin\mathscr{E}_{1},\ \{y,a_{0}\}\in\mathscr{E}_{1},\ \{x,y,a_{0}\}\in\mathscr{E}_{1},\ \ \$$\{x,a_{0}\}\notin\mathscr{E}_{2},\ \{y,a_{0}\}\in\mathscr{E}_{2},\ \{x,y,a_{0}\}\notin\mathscr{E}_{2}.$
For the choice on $D=\{x,y,a_{0}\}$ when we mix $\rho^{\succ_{1}}_{\mathscr{E}_{1}}$ and $\rho^{\succ_{2}}_{\mathscr{E}_{1}}$, $a_{0}$ is very good when the agent has preference $\succ_{2}$ but very bad when the agent has preference $\succ_{1}$. However, our model does not have the correlation between preferences and $\lambda$. Thus in order to rationalize this choice function (the mixture), in particular, the correlation between the preference relation and the value of the outside option, we need to construct preferences in a way that mimic the correlation pattern as follows. Each vertex gets a special underlying alternative $\overline{x}_{i}\in X(a_{0})$ that plays the role of the outside option only for $\succ_{i}$. That is, $\overline{x}_{i}$ is ranked at the top of $\succ_{i}$ when extended to $\mathscr{X}$. In particular, we can make $\overline{x}_{i}$ ranked very low in any ranking except $\succ_{i}$ (when extended to $\mathscr{X}$). With this construction, $\lambda$ has a different effect on each $\succ_{i}$, which mimics the effect of correlation between $\mu_{{\scriptscriptstyle\mathscr{X}}}$ and $\lambda$. The 1-1 comes from the fact that we need another underlying alternative in $X(a_{0})$ at the bottom in that case where $a_{0}$ is never chosen.
Remark D.6.
Theorem 3.2 does not require the domain assumption that $a_{0}\in D$ for all $D\in\mathscr{D}$. The theorem continues to hold with a slight modification of the definition of $c^{\succ}_{\mathscr{E}}$. The adjustment is needed because a set $D$ may fail to contain $a_{0}$ even when $D\notin\mathscr{E}$. The generalized definition is:
Remark D.7.
Some readers may think that assuming additive separable preferences over $\mathscr{X}$ may help to justify the use of ARUM. For example, Eales and Unnevehr (1982) suggests that aggregates should be defined to satisfy separability. This is not the case, as separability does not address unknown composition in a discrete choice setting.
To see this, notice that the non-overlappingness of preferences is equivalent to the condition that any preferences can be represented by the following form of utility function. Assuming $\mathscr{A}=\{a_{0},a_{1}\}$, $\mathscr{A}_{N}=\{a_{0}\}$, $\mathscr{X}=\{x_{1},x_{2},y\}$ for simplicity,
$w(x_{1}+x_{2},y)$
where $X(a_{0})=\{x_{1},x_{2}\}$ and $X(a_{1})=\{y\}$. To see why, refer to Theorem 4.1, which says that non-overlapping preferences are equivalent to preferences over aggregates. Since $x_{1}+x_{2}$ is exactly the quantity of the outside option consumed, $w$ is a utility function over aggregates.
Thus, additive separability of preferences is not enough to justify ARUM as the general form,
$w(u(x_{1},x_{2}),y),$
where $w$ is linear, is much weaker than non-overlapping preferences. For example, let $w$ be the sum, and $u(x_{1},x_{2})=2x_{1}+\frac{1}{2}x_{2}$. This induces the preference $x_{1}\succ y\succ x_{2}$. [^35] We abuse notation and conflate the quantity with the good. Suppose that alternative $x$ in a choice set $D_{1}$ means $x_{1}$; alternative $x$ in choice set $D_{2}$ means $x_{2}$. Then we observe a cycle that $x\succ y$, which is revealed from a choice from $D_{1}$, and $y\succ x$, which is revealed from a choice from $D_{2}$. Thus there is no linear order (or utility function) over aggregates.
Remark D.8.
Proposition 4.1 says that ARUM is equivalent to rationalization by non-overlapping preferences (even if the composition distribution is menu dependent) and Proposition 4.2 says that ARUM is equivalent to menu independence (even if preferences are overlapping). These are two distinct, but not contradictory characterizations. Because because $(\mu_{\mathscr{X}},\lambda)$ is not identified, it is possible for an ARUM is generated by overlapping preferences and/or menu-dependent composition distribution. The two propositions respectively say that there exists a non-overlapping rationalization and a menu-independent rationalization.
Remark D.9.
To better understand the structure of the RU-rational set, in particular the non-convexity, note that $\rho$ is bilinear but not linear in $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$.
To see this, denote $\rho$ rationalized with $(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)$ by $\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}$. Fix a composition correspondence $X$. With $\mu_{{\scriptscriptstyle\mathscr{X}}}$ fixed, $\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}$ is linear in $\lambda$. For any $\alpha\in[0,1]$, composition distributions $\lambda$, $\lambda^{\prime}$ such that $\lambda^{\prime\prime}=\alpha\lambda+(1-\alpha)\lambda^{\prime}$, we have $\alpha\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}+(1-\alpha)\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda^{\prime})}=\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda^{\prime\prime})}$. Similarly, with $\lambda$ fixed, $\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}$ is linear in $\mu_{{\scriptscriptstyle\mathscr{X}}}$ in the sense that if $\alpha\mu_{{\scriptscriptstyle\mathscr{X}}}+(1-\alpha)\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime}=\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime\prime}$ then $\alpha\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}+(1-\alpha)\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}=\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime\prime},\lambda)}$.
However, it is not linear because $\alpha\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}+(1-\alpha)\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime},\lambda^{\prime})}\neq\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}}^{\prime\prime},\lambda^{\prime\prime})}$ in general. Note that this non-linearity of $\rho_{(\mu_{{\scriptscriptstyle\mathscr{X}}},\lambda)}$ does not directly implies the non-convexity of the RU rational set, although if it were linear then the set would immediately be convex.
Appendix E Supplementary material for simulation
E.1 Additional Explanation for Section 5
In this section, we provide further explanations of the figures in Section 5.
E.1.1 Figure 2b
E.1.2 Figure 5a
To understand the figure, recall that we consider three alternatives $x,y$, and $a_{0}$. Thus, it can be shown that $\rho$ is ARUM if and only if it satisfies the standard monotonicity condition for all menus. Since $\rho$ is RU-rational, it satisfies Limited Monotonicity, that is monotonicity on $\{x,y\}$ with respect to adding $a_{0}$. Also, since $\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$, monotonicity is satisfied on menu $\{y,a_{0}\}$ with respect to adding $x$. Thus the only possible ways to violate monotonicity are (i) $\rho(\{x,y,a_{0}\},a_{0})>\rho(\{x,a_{0}\},a_{0})$ and (ii) $\rho(\{x,y,a_{0}\},a_{0})>\rho(\{y,a_{0}\},a_{0})$. The first occurs when $a_{0}$ is less attractive on $\{x,a_{0}\}$ than on $\{x,y,a_{0}\}$ and the second occurs when $a_{0}$ is more attractive on $\{x,a_{0}\}$ than on $\{x,y,a_{0}\}$. Since $u(z)$ is high and $u(w)$ is low, the first case occurs when $a_{0}$ is more likely to be $\{w\}$ (e.g., $\lambda_{\{x,a_{0}\}}(\{w\})=1$); the second case occurs when $a_{0}$ is more likely to be $\{z\}$ or $\{z,w\}$ (e.g., $\lambda_{\{x,a_{0}\}}(\{w\})=0$). These are two cases where the distance is large in Figure 5a.
E.1.3 Figure 5b
In the following, we provide the results of simulations using different parameter values.
E.2 Estimation bias
E.2.1 Effect of composition distributions $\lambda$